OpenStudy (anonymous):

graph y=5/2cos(x-pi/4) and plot x-intercepts, maxima, minima for one period

OpenStudy (mosaic):

To find the x-intercepts: When does the cosine function become zero?

OpenStudy (anonymous):

set y=0?

OpenStudy (mosaic):

Well, in one period, y = cos(x) becomes zero when x = pi/2 and 3pi/2 Therefore, y=5/2cos(x-pi/4) becomes zero when x - pi/4 = pi/2 and x - pi/4 = 3pi/2 or x = pi/2 + pi/4 or x = 3pi/2 + pi/4. x = 3pi/4 or x = 7pi/4. These are the x-intercepts. At what x does y = cos(x) attain the maximum and minimum?

OpenStudy (anonymous):

the maximum between 3pi/4 and 7pi/4. and the minimum before 3pi/4?

OpenStudy (mosaic):

y = cos(x) attains its maximum at x = 0 and minimum at x = pi. Therefore, y = 5/2cos(x-pi/4) attains its maximum at x-pi/4 = 0 and minimum at x - pi/4 = pi. Solve for x. The max value (that is, the y value) will be +5/2. The min value will be -5/2.

OpenStudy (anonymous):

so if i plotted the min and max on a graph would it be (0, 5/2) for the max and (pi, -5/2) for the min? thank you so much by the way :)

OpenStudy (mosaic):

No. y = cos(x) attains it max at x = 0 and min at x = pi. y = 5/2cos(x-pi/4) attains its maximum when x-pi/4 = 0 and attains minimum when x - pi/4 = pi. You need to solve for x. max at x = pi/4 and min at x = pi + pi/4 = 5pi/4. Max: (pi/4, 5/2) Min: (5pi/4, -5/2)

OpenStudy (anonymous):

and this would all be within the period [0, 2pi]?

OpenStudy (mosaic):

Yes. pi/4 = 0.25pi which is in [0, 2pi] 5pi/4 = 1.25pi which is in [0, 2pi]

OpenStudy (mosaic):

0 < 0.25pi < 2pi 0 < 1.25pi < 2pi

OpenStudy (anonymous):

perfect. would you mind helping me graph it? i'm still a little confused on how it should be graphed?

OpenStudy (mosaic):

If you know the graph of y = cos(x), you can easily graph y = 5/2cos(x-pi/4). The first graph goes from -1 to +1. The second graph goes from -5/2 to +5/2. Since the second graph has x - pi/4, it would be like shifting the first graph to the RIGHT by pi/4.

OpenStudy (anonymous):

thank you. where would pi and 2pi be on the graph?

OpenStudy (mosaic):